OpenStudy (anonymous):

Simplify. -7(-4)^2+3(-2)2

OpenStudy (anonymous):

Start by multiplying out those squares.

OpenStudy (anonymous):

Expression is equivalent to -7*-4*-4+3*-2*-2 = -7*16+3*4 = -112+12 = -100